OpenStudy (anonymous):

Can someone help me solve 2x+4 = -3x+14

OpenStudy (anonymous):

The answer is x=2.

OpenStudy (anonymous):

thank you but I want to learn how to it I want to learn step by step if u know what I mean

OpenStudy (anonymous):

2x+3x = 14-4 5x = 10 x= 10/5 x = 2

OpenStudy (anonymous):

K, Iffat0311 had it right.

OpenStudy (anonymous):

thank you guys :)

OpenStudy (anonymous):

Anytime! Studying makes perfect

OpenStudy (anonymous):

i wish i could give everyone medal

OpenStudy (anonymous):

=)